Network Troubleshooting

Troubleshooting Tools

         Nick Urbanik <nicku(at)nicku.org>
Copyright Conditions: GNU FDL (see http://www.gnu.org/licenses/fdl.html)

References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 2 Introduction

Focus: Basics and Standard Tools . . . . . . . . . . . . . . . . . . . . slide 3 Troubleshooting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 4 Troubleshooting: Learn as you go . . . . . . . . . . . . . . . . . . . . slide 5

Dcoumentation Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 6 Documentation Tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 7

General Troubleshooting Problem Solving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 9 Identify the Problem . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 10 Gather the Facts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 11 Consider Possibilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 12 Action Plan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 13 Implement Action Plan . . . . . . . . . . . . . . . . . . . . . . . . . . slide 14 Observe Results. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 15 document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 16 modify action plan. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 17

TCP/IP OSI—TCP/IP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 19 IP Header—Layer 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 20 IP Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 21 IP Header (continued) . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 22 TCP Header—Layer 4 . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 23 TCP Header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 24 TCP Header (continued) . . . . . . . . . . . . . . . . . . . . . . . . . slide 25 UDP Header—Layer 4 . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 26

Troubleshooting TCP/IP Troubleshooting TCP/IP . . . . . . . . . . . . . . . . . . . . . . . . . slide 28 Troubleshooting TCP/IP . . . . . . . . . . . . . . . . . . . . . . . . . slide 29

Host Configuration Host Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 30 Host Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 31 Boot Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 32 Boot Windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 33 Determine Addresses . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 34 MAC ↔ IP mapping — 1. . . . . . . . . . . . . . . . . . . . . . . . . slide 35 MAC ↔ IP mapping — 2. . . . . . . . . . . . . . . . . . . . . . . . . slide 36 Routing Table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 37 Access Controls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 38 DNS resolver. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 39 Checking services provided . . . . . . . . . . . . . . . . . . . . . . . . slide 40 Server Running? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 41 top. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 42 netstat -tua . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 43 lsof: List Open Files . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 44 ifconfig . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 45 route . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 46

Cables

Connectivity Testing: Cabling . . . . . . . . . . . . . . . . . . . . . slide 47 Ping

Software tools: ping . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 49 Good ping, Bad ping?. . . . . . . . . . . . . . . . . . . . . . . . . . . slide 50 How to Use ping? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 51 fping: flood ping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 52 hping2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 53 arping: uses ARP requests . . . . . . . . . . . . . . . . . . . . . . . slide 54

traceroute

Path Discovery: traceroute . . . . . . . . . . . . . . . . . . . . . . slide 55 traceroute: How it Works . . . . . . . . . . . . . . . . . . . . . . . slide 56 traceroute Limitations . . . . . . . . . . . . . . . . . . . . . . . . . . slide 57

Measurements Performance Measurements: delay. . . . . . . . . . . . . . . . . . . slide 58 bandwidth and throughput. . . . . . . . . . . . . . . . . . . . . . . . slide 59 Quality of a Link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 60 Throughput: ping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 61 Througput: ping One . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 62 Throughput ping 2 remote. . . . . . . . . . . . . . . . . . . . . . . . slide 63 Throughput ping 2 remote. . . . . . . . . . . . . . . . . . . . . . . . slide 64 Throughput ping 2 remote. . . . . . . . . . . . . . . . . . . . . . . . slide 65 Throughput ping 2 remote. . . . . . . . . . . . . . . . . . . . . . . . slide 66 Throughput ping 2 remote. . . . . . . . . . . . . . . . . . . . . . . . slide 67 Limitations of ping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 68 Path Performance: Other tools . . . . . . . . . . . . . . . . . . . . . slide 69 pathchar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 70 Measuring Throughput. . . . . . . . . . . . . . . . . . . . . . . . . . . slide 71 Measuring Throughput with ttcp . . . . . . . . . . . . . . . . . . . slide 72

iproute The ip program, iproute . . . . . . . . . . . . . . . . . . . . . . . . . slide 74 iproute and iptables . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 75 Traffic Measurements: netstat -i . . . . . . . . . . . . . . . . . . slide 76 Measuring Traffic: netstat -i . . . . . . . . . . . . . . . . . . . . . slide 77 Traffic measurements: ifconfig, ip . . . . . . . . . . . . . . . . . slide 78 Getting more info using ip . . . . . . . . . . . . . . . . . . . . . . . . slide 79 Guide to ip 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 80 Guide to ip 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 81

Packet Capture What is Packet Capture? . . . . . . . . . . . . . . . . . . . . . . . . . slide 83 When Packet Capture? . . . . . . . . . . . . . . . . . . . . . . . . . . slide 84 Warning: Don’t Get Sacked!. . . . . . . . . . . . . . . . . . . . . . . slide 85 tcpdump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 86 How to use tcpdump. . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 87 tcpdump: some options. . . . . . . . . . . . . . . . . . . . . . . . . . . slide 88 tcpdump Filters: host and port . . . . . . . . . . . . . . . . . . . . . slide 89 tcpdump filters: networks . . . . . . . . . . . . . . . . . . . . . . . . . slide 90 tcpdump filters: protocol . . . . . . . . . . . . . . . . . . . . . . . . . slide 91 tcpdump filters: combining . . . . . . . . . . . . . . . . . . . . . . . . slide 92 Writing data to a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 93 Reading a Dumped File . . . . . . . . . . . . . . . . . . . . . . . . . . slide 94 HTTP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 95 tcpdump: When reading TCP . . . . . . . . . . . . . . . . . . . . . . slide 96 Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 97

Ethereal Ethereal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 99 Ethereal Screenshot . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 100 Ethereal Screenshot . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 101 You can expand any protocol:. . . . . . . . . . . . . . . . . . . . . slide 102 Ethereal 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 103 Display Filters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 104 Tools → Follow TCP Stream . . . . . . . . . . . . . . . . . . . . . slide 105 Ntop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 106

Ntop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 107 Switched Networks

Port Monitoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 109 How monitor one machine?. . . . . . . . . . . . . . . . . . . . . . . slide 110 Are switched networks secure? . . . . . . . . . . . . . . . . . . . . slide 111

Port Scanning What is a port scanner? . . . . . . . . . . . . . . . . . . . . . . . . . slide 113 How does nmap identify OS? . . . . . . . . . . . . . . . . . . . . . . slide 114 Running nmap: Use xnmap . . . . . . . . . . . . . . . . . . . . . . . slide 115 Uses of nmap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 116

DNS troubleshooting DNS troubleshooting . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 118 DNS: dig . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 119 dig: Checking forward DNS lookup. . . . . . . . . . . . . . . . . slide 120 dig: reverse lookup 1. . . . . . . . . . . . . . . . . . . . . . . . . . . slide 121 dig syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 122 dig: axfr (Zone Transfer) . . . . . . . . . . . . . . . . . . . . . . . . slide 123 nslookup: an interactive program . . . . . . . . . . . . . . . . . . slide 124 nslookup: reverse lookups . . . . . . . . . . . . . . . . . . . . . . . slide 125

telnet Email: testing with telnet . . . . . . . . . . . . . . . . . . . . . . slide 127 Test the VTC mail server: . . . . . . . . . . . . . . . . . . . . . . . slide 128 SMTP commands for sending mail . . . . . . . . . . . . . . . . . slide 129 Testing the VTC pop3 server 1 . . . . . . . . . . . . . . . . . . . . slide 130 Testing the pop3 server 2 . . . . . . . . . . . . . . . . . . . . . . . . slide 131 pop3 commands: retrieving mail . . . . . . . . . . . . . . . . . . . slide 132 telnet: Testing Other Applications . . . . . . . . . . . . . . . . slide 133 Conclusion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 134

References

• Joseph D. Sloan, Network Troubleshooting Tools, O’Reilly, August 2001, ISBN: 059600186-X, TK5105.5 .S557 2001

• Cisco’s  Troubleshooting  Overview  at:

   http://www.cisco.com/univercd/cc/td/doc/cisintwk/itg_v1/tr1901.htm

• Craig Hunt, TCP/IP Network Administration, Third Edition O’Reilly, April 2002, ISBN: 059600297-1. Chapter 13 covers troubleshooting TCP/IP.

• Martin A. Brown, Guide to IP Layer Network Administration with

  Linux, http://linux-ip.net/

• Noah Davids, Don’t forget to check your ARP cache,

  http://members.cox.net/~ndav1/self_published/The_ARP_cache.doc

SNM — ver. 1.4 Network Troubleshooting — slide 2

Focus: Basics and Standard Tools

• Solving network problems depends a lot on your understanding

• Simple tools can tell you what you need to know

• Example: ping is incredibly useful!

SNM — ver. 1.4 Network Troubleshooting — slide 3

Troubleshooting

• Avoid it by:

∘ redundancy ∘ documentation ∘ training

• Try quick fixes first ∘ simple problems often have big effects: ∘ is the power on? ∘ is the network cable plugged into the right socket? Is LED

flashing? ∘ has anything changed recently?

• Change only one thing at a time ∘ test thoroughly after the change

• Be familiar with the system ∘ maintain documentation

• Be familiar with your tools

∘ before trouble strikes

SNM — ver. 1.4 Network Troubleshooting — slide 4

Troubleshooting: Learn as you go

• Study and be familiar with the normal behaviour of your network • Monitoring tools can tell you when things are wrong

∘ if you know what things look like when they are right

• Using tools such as Ethereal can help you understand ∘ your network, and ∘ tcp^{/}ip— better

SNM — ver. 1.4 Network Troubleshooting — slide 5

Documentation

• Maintain an inventory of equipment and software ∘ a list mapping mac addresses to machines can be very helpful

• Maintain a change log for each major system, recording: ∘ each significant change ∘ each problem with the system ∘ each entry dated, with name of person who made the entry

• Two categories of documentation:

∘ Configuration information – describes the system

– use system tools to obtain a snapshot, e.g., sysreport in

Red Hat Linux ∘ Procedural information

– How to do things

– use tools that automatically document what you are do-

ing, e.g., script

SNM — ver. 1.4 Network Troubleshooting — slide 6

Documentation Tools

• Use script:
$ script ∼/logs/logfile-$(date +%F-%R).log

∘ starts a new shell ∘ all you type, all output goes into the file ∘ Add comments with # I tried this...

• Use tee:

$ arp -a | tee outfile

• Use sudo: all commands are recorded in /var/log/secure

• Use plod from http://bullwinkle.deer-run.com/~hal/plod/

∘ lets you record a worksheet easily ∘ Perl, so fine on any platform

SNM — ver. 1.4 Network Troubleshooting — slide 7

General Troubleshooting

Problem Solving

SNM — ver. 1.4 Network Troubleshooting — slide 9

Identify the Problem

• Problem is reported by a person or by software

• Often involves communicating with others ∘ Somewhat like gathering requirements in software design ∘ An iterative process

• Possible questions to ask:

∘ What does not work? ∘ What does work? ∘ Are the things that do and do not work related? ∘ Has the thing that does not work ever worked? ∘ When the problem was first noticed? ∘ What has changed since the last time it did work? ∘ Did anything unusual happen since the last time it worked? ∘ When exactly does the problem occur? ∘ Can the problem be reproduced and if so, how can it be re-

produced?

SNM — ver. 1.4 Network Troubleshooting — slide 10

Gather the Facts

• You probably need to find out more about the problem from other sources, including

∘ Asking other people: affected users, network administrators, managers, and other key people

∘ Network management systems, such as Nagios

  http://nagios.org/
∘ Tools such as Ethereal, tcpdump, ntop (http://ntop.org/)

— see slides starting at §83

∘ Server log files

∘ Documentation about your servers and network created by local staff

∘ Documentation about software and hardware that are pro-

vided by the vendors

SNM — ver. 1.4 Network Troubleshooting — slide 11

Consider Possibilities based on Facts

• Using the information you have gathered, try to eliminate some

potential problems from your list.

SNM — ver. 1.4 Network Troubleshooting — slide 12

Create an Action Plan

• Start with the most likely ∘ … and those that are easiest to test

• Plan needs to be methodical

• Plan to change only one thing at a time ∘ You can then understand the cause of the problem ∘ Aim to understand the problem so you can learn from it,

solve (or prevent) similar problems in the future

• Aim higher than just removing the symptoms!

SNM — ver. 1.4 Network Troubleshooting — slide 13

Implement Action Plan

• Perform each step carefully

• Test to see if symptoms go away

SNM — ver. 1.4 Network Troubleshooting — slide 14

Observe Results

• Gather results as you change each variable

• Use same techniques you used in slide 11:

∘ Check with the key people ∘ Check with your tools

SNM — ver. 1.4 Network Troubleshooting — slide 15

If Solved: Document Solution

• Record the problem and its resolution in the documentation you maintain for your company.

• Ensure others in your team can benefit from the insight you have

gained

SNM — ver. 1.4 Network Troubleshooting — slide 16

Otherwise, Modify Action Plan

• Go back to the steps in slide §13:

∘ Modify your action plan, selecting the next most likely action from your list

∘ You may have discovered more information in your investigation, so this can help you focus on likely causes.

• If you have exhausted all the items on your list, and cannot think of what else to do:

∘ Get help from the vendor

∘ Get help from mailing lists

∘ Discuss the problem with your network of colleagues (e.g., the people who are now studying with you, but who move on to work in a similar field!)

∘ You could even track me down and ask me! Quite a few of

my ex-students do.

SNM — ver. 1.4 Network Troubleshooting — slide 17

TCP/IP

OSI—TCP/IP IP Header—Layer 3

SNM — ver. 1.4 Network Troubleshooting — slide 20

IP Header

• Version — this is a 4-bit IP header length field that indicates the version of IP currently used. The current version of IP is 4 (IPv4) but IPv6 is already being implemented experimentally and will be supported on future versions of the IOS.

• IP Header Length (IHL) — this indicates the datagram header length in 32-bit words.

• Type of Service (ToS) — ToS specifies how a particular upperlayer protocol would like the current datagram to be handled. Datagrams can be assigned various levels of importance with this field.

• Total length — this specifies the length of the entire IP packet, including data and header, in bytes.

• Identification — this field contains an integer that identifies the current datagram. This field is used to help piece together datagram fragments.

SNM — ver. 1.4 Network Troubleshooting — slide 21

IP Header (continued)

• Flags — a flag is a 3-bit field of which the 2 low-order bits control fragmentation. One bit specifies whether the packet can be fragmented; the second bit specifies whether the packet is the last fragment in a series of fragmented packets.

• Time-to-Live — this field maintains a counter that gradually decrements down to zero, at which point the datagram is discarded. This prevents packets from looping endlessly.

• Protocol — protocol indicates which upper-layer protocol receives incoming packets after IP processing is complete.

• Header Checksum — this field helps ensure IP header integrity.

• Source Address — this field specifies the sending node.

• Destination Address — this field specifies the receiving node.

• Options — this field allows IP to support various options, such as security.

• Data — this field contains upper-layer information.

SNM — ver. 1.4 Network Troubleshooting — slide 22

TCP Header—Layer 4

SNM — ver. 1.4 Network Troubleshooting — slide 23

TCP Header

• Source port and destination port — these fields identify the points at which upper-layer source and destination processes receive TCP services.

• Sequence number — this field usually specifies the number assigned to the first byte of data in the current message. Under certain circumstances, it can also be used to identify an initial sequence number to be used in the upcoming transmission.

• Acknowledgment number — this field contains the sequence number of the next byte of data the sender of the packet expects to receive.

• Data offset — this field indicates the number of 32-bit words in the TCP header.

• Reserved — this field is reserved for future use.

• Flags — this field carries a variety of control information.

SNM — ver. 1.4 Network Troubleshooting — slide 24

TCP Header (continued)

• Window — this field specifies the size of the sender’s receive window (that is, buffer space available for incoming data).

• Checksum — this field indicates whether the header was damaged in transit.

• Urgent pointer — this field points to the first urgent data byte in the packet.

• Options — this field specifies various TCP options.

• Data — this field contains upper-layer information.

SNM — ver. 1.4 Network Troubleshooting — slide 25

UDP Header—Layer 4

• Source and Destination Port fields serve the same functions as they do in the TCP header.

• Length field specifies the length of the UDP header and data • Checksum field allows packet integrity checking. It is optional.

SNM — ver. 1.4 Network Troubleshooting — slide 26

Troubleshooting TCP/IP

Troubleshooting TCP/IP

SNM — ver. 1.4 Network Troubleshooting — slide 28

Troubleshooting TCP/IP

Step 1 First, determine whether your local host is properly configured (for instance, correct subnet mask and default gateway configuration).

Step 2 Next, use the ping or traceroute commands to determine whether the routers through which you must communicate can respond. Start with the most local router and progressively ping outwards through the Internet or use traceroute.

Step 3 If you cannot get through a particular node, examine the node configuration and use the various show commands to determine the state of the router (these include show ip route, show ip

arp, show running-configuration, and so on.)

Step 4 If you can get to all the routers in the path, check the host configuration at the remote host (or get someone’s help to do so), and check its configuration.

SNM — ver. 1.4 Network Troubleshooting — slide 29

Host Network Configuration tools

ps — information about processes top — dynamic information about processes netstat — show connections and services, routing lsof — list open files ifconfig — shows and changes network interfaces route — shows, changes routing table ip — show, change, set network configuration arp — shows MAC addresses nmap — portscanner: shows open ports, identifies os

SNM — ver. 1.4 Network Troubleshooting — slide 30

Checking (and Setting) Host Configuration

• Solving Boot problems: §32, §33

• Determine ip address, netmask, broadcast address: §34

• Deterine correct mac ↔ ip address mapping: §35, §36

• Examine routing table: §37

• Examine access controls: §38

• Examine web proxy settings: check web browser

• Examine dns resolver settings: §39

• Determine services provided: §40, §41

• Determine cpu, memory load conditions (is the server over-

loaded?) §42

SNM — ver. 1.4 Network Troubleshooting — slide 31

Boot problems: Linux

• Use grub to interactively boot the computer (see my extensive

  grub handout: http://nicku.org/ossi/lab/grub/grub.pdf)

• Verify that /etc/fstab mounts the correct filesystems

• Use a rescue disk such as Knoppix or the Red Hat installation cdrom.

• This gives you full access to the system and repairing boot prob-

lems is pretty straightforward.

SNM — ver. 1.4 Network Troubleshooting — slide 32

Boot problems: Windows

• Use the installation Windows cd to enter the (extremely limited) system repair mode. I believe this is called the recovery console.

• Use  the  Linux  floppy  bootdisk  at

  http://home.eunet.no/~pnordahl/ntpasswd/ to replace

the Administrator password

• Use  the  bootable  Windows  cdrom:

  http://www.nu2.nu/pebuilder/;

∘ Gives full access to the NTFS file system.

∘ Not as good with Windows as Knoppix is with Linux, but better than another reinstall.

∘ takes some time to build.

∘ Henry Leung (in A204d) has built some.

SNM — ver. 1.4 Network Troubleshooting — slide 33

Determine Addresses

Linux: On Linux, these commands all show the ip address, mac address, netmask and broadcast address for all (or the specified) interface. The commands ip and ifconfig are in the directory

/sbin; netstat is in /bin.
$ ip addr
$ ip addr show eth0
$ ifconfig
$ ifconfig eth0
$ netstat -i

Windows:

C:\> ipconfig /all

Cisco IOS: these are both privileged commands, as shown by the prompt:

Router# show running-config
Router# show ip interface brief
Router# show ip interface

SNM — ver. 1.4 Network Troubleshooting — slide 34

MAC ↔ IP mapping — 1

Linux:

     $ arp -a
     $ ip neigh show

The lifetime of the arp cache entries is settable in /proc/sys/

     net/ipv4/neigh/⟨interface⟩/gc stale time and is normally 60

seconds.

Cisco IOS:

     Router# show ip arp
Note that this document: http://members.cox.net/~ndav1/self_published/The_ARP_cache.doc

has a good discussion on troubleshooting arp. The online book at http://linux-ip.net/ has an excellent chapter on arp. arp is probably the most dangerously exposed protocol in a lan, and can easily be subverted by tools such as dsniff and Ettercap. Hardcode important arp entries to avoid attack.

SNM — ver. 1.4 Network Troubleshooting — slide 35

MAC ↔ IP mapping — 2

Windows:

C:\> arp -a

You may wish to clear the arp cache on a Windows machine with: C:\> arp -d ⟨IP address⟩ or clear the entire arp cache with:

C:\> arp -d *

since the Windows arp cache lives 10 minutes by default, a rather (too?) long time. It can be changed by two registry entries under

HKEY_LOCAL_MACHINE\SYSTEM\CurrentControlSet\Services\
Tcpip\Parameters.

SNM — ver. 1.4 Network Troubleshooting — slide 36

Routing Table

Linux: The commands ip and route are in /sbin, the command

netstat is in /bin.
$ ip route
$ route -n
$ netstat -nr

Windows:

C:\> route print
C:\> netstat -nr

Cisco IOS:

Router# show ip route

SNM — ver. 1.4 Network Troubleshooting — slide 37

Access Controls

• Access controls can block access mysteriously unless you think to check for it.

Linux: There are two main things to check. The iptables command is in the /sbin directory.

$ iptables -L -n

Note that Linux and many other posix systems implement the tcpwrappers access control in /etc/hosts.allow and /etc/

hosts.deny. See man hosts.allow and man hosts.deny.

Cisco IOS:

Router# show ip access-list

SNM — ver. 1.4 Network Troubleshooting — slide 38

DNS resolver settings

Linux: The configuration for the resolver is /etc/resolv.conf. This determines what name servers the system will ask. It also tells what domain will be appended to a hostname.

The /etc/hosts file is usually the first way hostname ↔ ip address mappings are made, but this can be changed in /etc/ nsswitch.conf. To ask the operating system for what it can see there, do:

$ getent hosts

Linux provides three tools for troubleshooting dns and dns

servers: dig, host and nslookup.

Windows: See the output of

C:\> ipconfig /all

for the names of the dns server the resolver will use. Recent versions of Windows provide the program nslookup to help with dns troubleshooting.

See Slides starting at §118 for details of dig and nslookup.

SNM — ver. 1.4 Network Troubleshooting — slide 39

Checking services provided

Linux: There are four main ways to check: • Verify the processes with ps (see §41)

• Verify the services that are configured to start when the system boots:

$ chkconfig --list | grep on

• Check that the service is listening on the network interface:

$ netstat -tua

will show all network connections to this machine.

• The lsof program can be helpful in diagnosing problems

with network services. See §44.

Windows: Check network connections with

C:\> netstat -a

SNM — ver. 1.4 Network Troubleshooting — slide 40

Using ps to See If Server Running

• Is the network service running on the server?

• Is the web server running?

$ ps aux | grep httpd

• Is the DHCP server running?

$ ps aux | grep dhcpd

• Is the directory server running?

$ ps aux | grep slapd

• Windows: use the task manager

SNM — ver. 1.4 Network Troubleshooting — slide 41

Using top to see Resource Hogs

• The program top shows:

∘ load average (the average number of processes that are ready to run, but for which no cpu is available)

– a load average of 4 or more is “quite high” ∘ processes that use the most resources

SNM — ver. 1.4 Network Troubleshooting — slide 42

netstat -tua: See Network Connections

• netstat -tua shows all network connections, including those listening

• sudo netstat -tuap shows all network connections, including those listening, and the processes responsible

• netstat -tul shows all network listeners

• netstat -t shows only tcp connections that are established • netstat -i is like ifconfig, shows info and stats about each

interface

• netstat -nr shows the routing table, like route -n

• Windows provides netstat also.

SNM — ver. 1.4 Network Troubleshooting — slide 43

lsof: List Open Files

• An amazingly useful tool

• Available for almost any Unix system

• lsof -i shows output to Internet and X.25 files, but won’t show connections that have terminated

• lsof -i@nicku.org will show only connections to that machine • Can monitor progress of an FTP transfer, many, many other ap-

plications

• See manpage, faq and quick start guide.

• Apparently, no equivalent tool available on Windows.

SNM — ver. 1.4 Network Troubleshooting — slide 44

ifconfig

• ifconfig eth0 — show stats on network interface eth0

• sudo ifconfig lo 127.0.0.1 — configure the loopback interface, start it up

• sudo ifconfig eth0 172.19.233.5 netmask 255.255.255.0

— configure eth0 with IP address 172.19.233.5/24

• ifconfig — show all configured network interfaces

• ifconfig -a — show all interfaces, including those not configured

yet.

SNM — ver. 1.4 Network Troubleshooting — slide 45

route

• route -n — print routing table

• route add 127.0.0.1 — add a route to localhost;

• should have been done automatically when created device with

  ifconfig
• route add -net 172.19.233.0 — add a route to the eth0 con-

figured on previous slide

• should have been done automatically by ifconfig

• route add 172.19.64.0 gw 172.19.233.254 — add a static

route to network 172.19.64.0 through router 172.19.233.254

• route add default gw 172.19.233.253 — add a default route

to 172.19.233.253 through eth0

SNM — ver. 1.4 Network Troubleshooting — slide 46

Connectivity Testing: Cabling

• Label cables clearly at each end

• Cable testers ∘ ensure wired correctly, check: ∘ attenuation ∘ length — is it too long?

– 100BaseT: less than 100m • Is the activity light on the interface blinking?

SNM — ver. 1.4 Network Troubleshooting — slide 47

Ping

Software tools: ping

• Most useful check of connectivity

• Universal

• If ping hostname, includes a rough check of dns

• Sends an icmp (Internet Control Message Protocol) ECHO REQUEST • Waits for an icmp ECHO REPLY

• Most pings can display round trip time

• Most pings can allow setting size of packet

• Can use to make a crude measurement of throughput—see §61

SNM — ver. 1.4 Network Troubleshooting — slide 49

What ping Result is Good, Bad?

• A steady stream of consistent replies indicates probably okay • Usually first reply takes longer due to arp lookups at each router

∘ After that, arp results are cached

• icmp error messages can help understand results:

∘ Destination Network Unreachable indicates the host doing

ping cannot reach the network ∘ Destination Host Unreachable may come from routers further

away

SNM — ver. 1.4 Network Troubleshooting — slide 50

How to Use ping?

• Ensure local host networking is enabled first: ping localhost, local ip address

• ping a known host on local network

• ping local and remote interfaces on router

• ping by IP as well as by hostname if hostname ping fails ∘ confirm DNS with dig (or nslookup) — see slide §118

• Ping from more than one host

SNM — ver. 1.4 Network Troubleshooting — slide 51

fping: flood ping

• Designed to test a large number of hosts

• more efficient than ping

• Used extensively by monitoring software such as

  mon:  http://www.kernel.org/software/mon/, nagios:
  http://www.nagios.org/

• take care not to flood too much!

• rpms are available; I built one (a long time ago) and put it on

ictlab under ∼ftp/pub/redhat/contrib

SNM — ver. 1.4 Network Troubleshooting — slide 52

hping2: ping anything with anything

• able to send custom TCP/IP packets and

• display target replies like ping program does with icmp replies.

• Can install with

  $ yum -y install hping2

on Fedora Core 1.

• See http://www.hping.org/.

SNM — ver. 1.4 Network Troubleshooting — slide 53

arping: uses ARP requests

• Limited to local network

• Can work with mac or ip addresses

• use to probe for arp entries in router (very useful!)

• packet filtering ∘ can block icmp pings, but ∘ won’t block arp requests

SNM — ver. 1.4 Network Troubleshooting — slide 54

Path Discovery: traceroute

• Sends udp packets ∘ (Microsoft tracert sends icmp packets)

• increments Time to Live (ttl) in ip packet header

• Sends three packets at each ttl

• records round trip time for each

• increases ttl until enough to reach destination

SNM — ver. 1.4 Network Troubleshooting — slide 55

traceroute: How it Works

• As ip packets pass through each router, ttl in ip header is decremented

• Packet is discarded when TTL decrements to 0

• ROUTER sends icmp TIME EXCEEDED message back to traceroute host

• When udp packet reaches destination, gets icmp

PORT UNREACHABLE, since uses an unused high udp port

SNM — ver. 1.4 Network Troubleshooting — slide 56

traceroute Limitations

• Each router has a number of ip addresses

• but traceroute only shows the one it used

• get different addresses when run traceroute from other end • sometimes route is asymmetric

• router may be configured to not send icmp TIME EXCEEDED mes-

sages

∘ get stars: * instead of round-trip time in traceroute output

SNM — ver. 1.4 Network Troubleshooting — slide 57

Performance Measurements: delay

• Three sources of delay:

• transmission delay — time to put signal onto cable or media ∘ depends on transmission rate and size of frame

• propagation delay — time for signal to travel across the media ∘ determined by type of media and distance

• queuing delay — time spent waiting for retransmission in a router

SNM — ver. 1.4 Network Troubleshooting — slide 58

Is Bandwidth == Throughput?

• bandwidth — the difference between the upper frequency and the

lower frequency that a channel can carry

∘ measured in Hertz

• throughput — amount of data that can be sent over link in given

time

∘ is not the same as bandwidth, which really has no direct

meaning with digital information

• bandwidth is related to throughput by the Shannon-Hartley Theorem; throughput ∝ bandwidth if signal to noise ratio is fixed:

C_{max} = B log_{2} (1 + \frac{S}{N} ) bits/sec
where C_{max} = maximum channel capacity,
B = bandwidth in Hz,
S = signal power in W,
N = noise power in W.

SNM — ver. 1.4 Network Troubleshooting — slide 59

Quality of a Link

• Other measurements needed

∘ i.e., for quality of service for multimedia

SNM — ver. 1.4 Network Troubleshooting — slide 60

ping Roughly Estimating Throughput

• Example: measuring throughput between this machine and one remote machine.

• ping with packet size = 100 bytes, round-trip time = 30ms

• ping with packet size = 1100 bytes, round-trip time = 60ms

• So takes 30ms extra (15ms one way) to send additional 1000 bytes, or 8000 bits

• Throughput is roughly 8000 bits per 15ms, or about 530,000 bits per second

• A very crude measurement:

∘ no account for other traffic, treats all links on path, there

and back, as one. ∘ Routers sometimes send packets onwards with much higher priority than with which they answer pings. See slide §68.

SNM — ver. 1.4 Network Troubleshooting — slide 61

Throughput: ping One Remote Host

• This can be expressed as a simple formula:

T P = 16 × \frac{P_{l} − P_{s}}{t_{l} − t_{s}} bits per second, where
P_{l} = size of large packet
P_{s} = size of small packet
t_{l} = round-trip time for large packet
t_{s} = round-trip time for small packet

Here we have:

T P = 16 × \frac{1100 − 100}{(60 − 30) × 10^{−3}}

= 16 × \frac{1000}{30 × 10^{−3}}

= \frac{16}{30} × 10^{6}

≈ 530,000 bps

SNM — ver. 1.4 Network Troubleshooting — slide 62

Throughput: ping Two Remote Hosts

• Measure throughput between two remote hosts: may use tools like

     ping

• ping two locations with two packet sizes (4 pings altogether, minimum)

∘ Many ping programs calculate average ping time: better to make a number of pings, use the average ping time.

∘ First ping time may be longer due to the time to get an answer to the arp request

∘ May be better to ping once, then start pinging again, and use the average ping time.

• Example:

Address  RTT 100 bytes RTT 1100 bytes
205.153.61.1 1.380 ms  5.805 ms
205.153.60.2 4.985 ms  12.823 ms
165.166.36.17 8.621 ms  26.713 ms

SNM — ver. 1.4 Network Troubleshooting — slide 63

Throughput: ping Two Remote Hosts — 2

Address  RTT 100 bytes RTT 1100 bytes
205.153.61.1 1.380 ms  5.805 ms
205.153.60.2 4.985 ms  12.823 ms
165.166.36.17 8.621 ms  26.713 ms

• Time difference / 2 (round trip time (RTT) → one way)

• Divide by size difference in bits: 8000 = 8 × (1100 − 100)

• Multiply by 1000 (ms → seconds)

• Mbs = bps/10^{6}

Near link Far link  Time difference Est. Through- put
205.153.61.1 205.153.60.2 3.413 ms  4.69 Mbps
205.153.60.2 165.166.36.17 10.254 ms 1.56 Mbps
3.413 ms = (12.823 − 4.985) − (5.805 − 1.380) ms
10.254 ms = (26.713 − 8.621) − (12.823 − 4.985) ms

SNM — ver. 1.4 Network Troubleshooting — slide 64

Throughput: ping Two Remote Hosts — 3

T P = 16 × \frac{P_{l} − P_{s}}{t_{fl} − t_{fs} − t_{nl} + t_{ns}} bits per second

where:

P_{l} = large packet size, bytes
P_{s} = small packet size, bytes
t_{nl} = ping time for larger packet to the near link, seconds
t_{ns} = ping time for smaller packet to the near link, seconds
t_{fl} = ping time for larger packet to the f ar link, seconds
t_{fs} = ping time for smaller packet to the f ar link, seconds

SNM — ver. 1.4 Network Troubleshooting — slide 65

Throughput: ping Two Remote Hosts — 4

P_{l} = large packet size, bytes = 1100 bytes
P_{s} = small packet size, bytes = 100 bytes
t_{nl} = ping time for larger packet to the near link, seconds

= 5.805 × 10^{−3} seconds

t_{ns} = ping time for smaller packet to the near link, seconds

= 1.380 × 10^{−3} seconds

t_{fl} = ping time for larger packet to the f ar link, seconds

= 12.823 × 10^{−3} seconds

t_{fs} = ping time for smaller packet to the f ar link, seconds

= 4.985 × 10^{−3} seconds

T P = 16 × \frac{P_{l} − P_{s}}{t_{fl} − t_{fs} − t_{nl} + t_{ns}} bits per second

= 16 × \frac{1100 − 100}{(12.823 − 4.985 − 5.805 + 1.380) × 10^{−3}} bits per second

= 16 × \frac{1000}{3.413 × 10^{−3}}

≈ 4,687,958 ≈ 4.69 Megabits per second

SNM — ver. 1.4 Network Troubleshooting — slide 66

Throughput: ping Two Remote Hosts — 5

P_{l} = large packet size, bytes = 1100 bytes
P_{s} = small packet size, bytes = 100 bytes
t_{nl} = ping time for larger packet to the near link, seconds

= 12.823 × 10^{−3} seconds

t_{ns} = ping time for smaller packet to the near link, seconds

= 4.985 × 10^{−3} seconds

t_{fl} = ping time for larger packet to the f ar link, seconds

= 26.713 × 10^{−3} seconds

t_{fs} = ping time for smaller packet to the f ar link, seconds

= 8.621 × 10^{−3} seconds

T P = 16 × \frac{P_{l} − P_{s}}{t_{fl} − t_{fs} − t_{nl} + t_{ns}} bits per second

= 16 × \frac{1100 − 100}{(26.713 − 8.621 − 12.823 + 4.985) × 10^{−3}} bits per second

= 16 × \frac{1000}{10.254 × 10^{−3}}

≈ 1,560,366 ≈ 1.56 Megabits per second

SNM — ver. 1.4 Network Troubleshooting — slide 67

Limitations of measuring with ping

• Most modern routers give high priority to routing

∘ Expecially switching routers, such as the Cisco 6509 in our Campus

∘ Many give much lower priority to answering pings

∘ The difference can be so great that the ping reply sometimes comes sooner from a more distant router, which according to our formula, indicates a negative throughput!

∘ Do not blindly apply this formula!

• Measurements may not match the kind of traffic created by the application you support

• The big advantages of these icmp measurements are:

∘ you do not need access to the machines, and

∘ you do not need to install any special software on them.

SNM — ver. 1.4 Network Troubleshooting — slide 68

Path Performance: Other tools

• Could use a tool like pathchar, bing, clink, pchar, or tmetric

that performs this calculation for you

• Use http://www.google.com/ to locate these tools

• pathchar is only available in binary form

• Others in source form, need compile with commands something like this:

  $ cd bing-1.1.3
  $ make
  $ sudo make install

SNM — ver. 1.4 Network Troubleshooting — slide 69

Path measurement with pathchar

$ sudo ./pathchar sina.com.hk
pathchar to sina.com.hk (202.85.139.140)
can't find path mtu - using 1500 bytes.
doing 32 probes at each of 45 sizes (64 to 1500 by 32)
0 localhost (127.0.0.1)
|   106 Mb/s,   293 us (698 us),  +q 1.18 ms (15.7 KB)
1 172.19.35.246 (172.19.35.246)
|    28 Mb/s,   488 us (2.10 ms)
2 192.168.83.2 (192.168.83.2)
3  * 1   448 798       2
|    20 Mb/s,   273 us (3.25 ms)
4 cw7204.vtc.edu.hk (202.40.210.220)
|   6.8 Mb/s,   521 us (6.04 ms)
5 210.176.123.37 (210.176.123.37)
|    52 Mb/s,   20 us (6.31 ms)
6 210.87.254.61 (210.87.254.61)
|   136 Mb/s,   116 us (6.63 ms)
7 g5-0-0.wttbr01.imsbiz.com (210.87.254.129)
|    33 Mb/s,   0.94 ms (8.88 ms),  +q 1.48 ms (6.10 KB) *6
8 iadvantage3-RGE.hkix.net (202.40.161.172)
|   164 Mb/s,   45 us (9.04 ms),  +q 1.74 ms (35.6 KB) *6
9 v005-m02.hk01.iadvantage.net (202.85.129.53)
|   ?? b/s,   -66 us (8.88 ms)
10 202.85.129.136 (202.85.129.136)
|   ?? b/s,   459 us (9.79 ms)
11 202.85.139.11 (202.85.139.11)
11 hops, rtt 6.18 ms (9.79 ms), bottleneck 6.8 Mb/s, pipe 9361 bytes

SNM — ver. 1.4 Network Troubleshooting — slide 70

Measuring Throughput

• May use ftp to transfer a large file, measure time

∘ tests whole path

∘ problem: affected by disk i^{/}o, xinetd

• May use a web browser and measure download time

∘ Problem: may be affected by caching in the web browser

∘ May be affected by caching in web proxies

• Better: measure using traffic similar to that created by the application.

SNM — ver. 1.4 Network Troubleshooting — slide 71

Measuring Throughput with ttcp

• Use ttcp, not affected by disk i^{/}o

• Consists of a client and server

• Need have installed at both ends

• Part of Red Hat Linux, Cisco ios

• Example: first, start receiver on ictlab:

$ ttcp -r -s
ttcp-r: buflen=8192, nbuf=2048, align=16384/0, port=5001  tcp
ttcp-r: socket
ttcp-r: accept from 172.19.32.30
ttcp-r: 16777216 bytes in 1.45 real seconds = 11285.88 KB/sec +++
ttcp-r: 9704 I/O calls, msec/call = 0.15, calls/sec = 6684.46
ttcp-r: 0.0user 0.2sys 0:01real 14% 0i+0d 0maxrss 0+2pf 0+0csw

• Second, start transmitter on nickpc:

$ ttcp -t -s ictlab
ttcp-t: buflen=8192, nbuf=2048, align=16384/0, port=5001  tcp  -> ictlab
ttcp-t: socket
ttcp-t: connect
ttcp-t: 16777216 bytes in 1.45 real seconds = 11335.64 KB/sec +++
ttcp-t: 2048 I/O calls, msec/call = 0.72, calls/sec = 1416.95
ttcp-t: 0.0user 0.0sys 0:01real 4% 0i+0d 0maxrss 0+2pf 0+0csw

SNM — ver. 1.4 Network Troubleshooting — slide 72

iproute

The ip program, iproute

• The ip program in the iproute package provides complete control over tcp^{/}ip networking in a Linux system

• Provides more networking control facilities than other tcp^{/}ip implementations

• Supports tunneling in many forms

• iproute documentation is in two manuals, one for ip routing, the

other for tunnelling

SNM — ver. 1.4 Network Troubleshooting — slide 74

iproute and iptables

• Between these software packages, you can:

∘ throttle bandwidth for certain computers ∘ throttle bandwidth to certain computers ∘ fairly share bandwidth ∘ protect your network from DoS attacks ∘ protect Internet from your customers ∘ multiplex many servers into one, for load balancing or for

high availability ∘ restrict access to your computers ∘ limit access of your users to other hosts ∘ do routing based on user id, mac address, source ip, port,

type of service, time of day or content

• See the Linux Advanced Routing and Traffic Control HOWTO at

http://tldp.org for details

SNM — ver. 1.4 Network Troubleshooting — slide 75

Traffic Measurements: netstat -i

• The netstat program can show statistics about network interfaces • Linux netstat shows lost packets in three categories:

∘ errors,

∘ drops (queue full: shouldn’t happen!)

∘ overruns (last data overwritten by new data before old data

was read: shouldn’t happen!) ∘ drops and overruns indicate faulty flow control — bad!

• These values are cumulative (since interface was up)

• Could put a load on interface to see current condition, with ping -l, to send large number of packets to destination

• See the difference in values

SNM — ver. 1.4 Network Troubleshooting — slide 76

Measuring Traffic: netstat -i

• Here we run netstat -i on ictlab:
  $ netstat -i
  Kernel Interface table
    Iface  MTU Met     RX-OK RX-ERR RX-DRP RX-OVR      TX-OK TX-ERR TX-DRP TX-OVR Flg
    eth0  1500   0 407027830      0      0      0 1603191764      0      0      3 BMRU
    lo   16436   0   2858402      0      0      0    2858402      0      0      0 LRU

• Notice that of the 1.6 billion bytes transmitted, there were 3 overuns.

• Next, blast the path you want to test with packets using ping -l

or the spray program, and measure again.

SNM — ver. 1.4 Network Troubleshooting — slide 77

Traffic measurements: ifconfig, ip

• ifconfig and ip give more information than netstat -i:

$ ifconfig eth0
eth0      Link encap:Ethernet  HWaddr 00:00:E2:35:AF:EE
inet addr:172.19.64.52  Bcast:172.19.127.255  Mask:255.255.192.0
IPX/Ethernet 802.2 addr:33001601:0000E235AFEE
UP BROADCAST RUNNING MULTICAST  MTU:1500  Metric:1
RX packets:407579600 errors:0 dropped:0 overruns:0 frame:0
TX packets:1605655688 errors:0 dropped:0 overruns:3 carrier:0
collisions:0 txqueuelen:100
RX bytes:3055300191 (2913.7 Mb)  TX bytes:2048217058 (1953.3 Mb)
Interrupt:18 Base address:0xd000
$ ip -s link list eth0
2: eth0: <BROADCAST,MULTICAST,UP> mtu 1500 qdisc pfifo_fast qlen 100
link/ether 00:00:e2:35:af:ee brd ff:ff:ff:ff:ff:ff
RX: bytes  packets  errors  dropped overrun mcast
3058362227 407610495 0       0       0       0
TX: bytes  packets  errors  dropped carrier collsns
2140511920 1605768150 0       0       0       0

SNM — ver. 1.4 Network Troubleshooting — slide 78

Getting more info using ip

• The -s (-statistics) option to ip provides statistics. Adding a second gives you even more:

$ ip -s -s link list eth0
2: eth0: <BROADCAST,MULTICAST,UP> mtu 1500 qdisc pfifo_fast qlen 100
 link/ether 00:00:e2:35:af:ee brd ff:ff:ff:ff:ff:ff
 RX: bytes  packets   errors  dropped overrun mcast
 3070792102 407726727  0       0       0       0
 RX errors: length    crc     frame   fifo    missed
            0          0       0       0       0
 TX: bytes  packets   errors  dropped carrier collsns
 2445799644 1606151878 0       0       0       0
 TX errors: aborted   fifo    window  heartbeat
            0          3       0       0

SNM — ver. 1.4 Network Troubleshooting — slide 79

Quick Guide to using ip: set up interface

• Here we set up a network interface and give it the IP address 192.168.0.1/24:

  $ ip link set dev eth1 up
  $ ip addr add 192.168.0.1/24 brd + dev eth1

• Two important points:

∘ If you do not specify the netmask, a netmask of /32 is assumed

∘ brd + means obtain broadcast address by setting the host

bits

SNM — ver. 1.4 Network Troubleshooting — slide 80

Quick Guide to using ip: set up routes

$ ip route add default dev eth1 via 192.168.0.254
$ ip route add 192.168.1.0/24 via 192.168.0.10

• The last adds a static route to another network

• the first adds the default route.

• You can omit the device if the network can be reached through a particular interface without any ambiguity

∘ I.e., ip is smart enough to figure out which network device

to use, though specifying it doesn’t hurt.

SNM — ver. 1.4 Network Troubleshooting — slide 81

Packet Capture tcpdump, Ethereal and Ntop

What is Packet Capture?

• Real time collection of data as it travels over networks

• Tools called:

∘ packet sniffers ∘ packet analysers ∘ protocol analysers, and sometimes even ∘ traffic monitors

SNM — ver. 1.4 Network Troubleshooting — slide 83

When Packet Capture?

• Most powerful technique

• When need to see what client and server are actually saying to each other

• When need to analyse type of traffic on network

• Requires understanding of network protocols to use effectively

SNM — ver. 1.4 Network Troubleshooting — slide 84

Warning: Don’t Get Sacked!

• Be sure that your boss agrees with you capturing packets on your company’s network

• People have been sacked for doing this without permission!

• Some have suffered long lawsuits and criminal records:

    ∘ See  http://www.stonehenge.com/merlyn/,  and
      http://www.lightlink.com/spacenka/fors/ for a fa-

mous example

• Do not invade the privacy of others

∘ Capturing passwords with insecure protocols such as telnet, ftp, http (that is not encrypted with tls) is very easy

– DON’T DO IT!

SNM — ver. 1.4 Network Troubleshooting — slide 85

tcpdump

• Available everywhere

• Windows: http://windump.polito.it/

• Syntax also used by other programs (such as Ethereal)

• Often it is the only tool available, so good to know

• Works by putting network interface into promiscuous mode

∘ normal Ethernet interface will ignore packets not addressed

to it ∘ in promiscuous mode, will examine all packets that arrive,

even those not addressed to it

SNM — ver. 1.4 Network Troubleshooting — slide 86

How to use tcpdump

• Can just type its name (as root):

     $ sudo tcpdump

• … but get a huge amount of data!

• Can restrict the data collected using a filter

• A filter may select addresses, protocols, port numbers, …

SNM — ver. 1.4 Network Troubleshooting — slide 87

tcpdump: some options

-c ⟨n⟩ capture a \underline{c}ount of ⟨n⟩ packets then stop -w ⟨file⟩ \underline{w}rite raw data to ⟨file⟩.

• Very useful — can filter and analyse this later with tcpdump, ethereal or other tools

• but you cannot see what you are capturing till later! -i ⟨interface⟩ collect from ⟨interface⟩ instead of lowest numbered net-

work \underline{i}nterface -s ⟨bytes⟩ collect no more than ⟨bytes⟩ of data from each packet in-

stead of default 68 bytes -e show link level info, e.g., \underline{E}thernet addresses -x gives a he\underline{x}adecimal dump of packets excluding link level data -X display ascii as well as hexadecimal if have -x option too Many more options: man tcpdump

SNM — ver. 1.4 Network Troubleshooting — slide 88

tcpdump Filters: host and port

• Show all network traffic to and from 192.168.0.1:

  $ tcpdump host 192.168.0.1

• Show packets to 192.168.0.1:

  $ tcpdump dst 192.168.0.1

• Show packets to port 68 on 192.168.0.1:

  $ tcpdump dst 192.168.0.1 and port 68

SNM — ver. 1.4 Network Troubleshooting — slide 89

tcpdump filters: networks

• Capture traffic to or from 205.153.60/24:

  $ tcpdump net 172.19.64/18

• can specify network as source or destination:

  $ tcpdump src net 205.153.60/24
  $ tcpdump dst net 172.19.64/18

SNM — ver. 1.4 Network Troubleshooting — slide 90

tcpdump filters: protocol

• tcpdump ip
• tcpdump tcp
• tcpdump ip proto ospf

• This will catch DNS name lookups, but not zone transfers (which use tcp):

• tcpdump udp port 53

SNM — ver. 1.4 Network Troubleshooting — slide 91

tcpdump filters: combining

• This will not work as you might expect:

   • tcpdump host ictlab and udp or arp

• Instead, need group with parentheses, and quote:

   • tcpdump "host ictlab and (udp or arp)"

• many more ways of filtering: man tcpdump

SNM — ver. 1.4 Network Troubleshooting — slide 92

Writing data to a file

$ sudo tcpdump -c 1000 -w ∼/tmp/tcpdump.pcap
tcpdump: listening on eth0
1014 packets received by filter
0 packets dropped by kernel

SNM — ver. 1.4 Network Troubleshooting — slide 93

Reading a Dumped File

$ tcpdump -nr ∼/tmp/tcpdump.pcap arp
22:32:41.751452 arp who-has 172.19.127.254 tell 172.19.127.29
22:32:41.863173 arp who-has 172.19.64.52 tell 172.19.64.63
22:32:41.863198 arp reply 172.19.64.52 is-at 0:0:e2:35:af:ee
22:32:42.082584 arp who-has 172.19.65.16 tell 172.19.125.229
22:32:43.113655 arp who-has 172.19.123.211 tell 172.19.65.2
22:32:44.635149 arp who-has 172.19.65.16 tell 172.19.127.106
22:32:44.874117 arp who-has 172.19.65.6 tell 172.19.126.174
22:32:45.147178 arp who-has 172.19.65.16 tell 172.19.126.240
22:32:45.209507 arp who-has 172.19.127.254 tell 172.19.125.127
22:32:45.212484 arp who-has 172.19.127.175 tell 172.19.125.127
22:32:45.239445 arp who-has 172.19.127.254 tell 172.19.125.212
22:32:45.455863 arp who-has 172.19.65.16 tell 172.19.126.194
22:32:45.540507 arp who-has 172.19.126.50 (44:30:54:59:43:4d)
 tell 172.19.65.10
22:32:45.562004 arp who-has 172.19.126.50 tell 172.19.65.2

SNM — ver. 1.4 Network Troubleshooting — slide 94

HTTP

$ tcpdump -nr ∼/tmp/tcpdump.pcap port http
22:43:32.633636 192.168.25.9.14075 > 172.19.64.52.http:
 S 1015952778:1015952778(0) win 6144 <mss 1460> (DF)
22:43:32.633693 172.19.64.52.http > 192.168.25.9.14075:
 S 1929920485:1929920485(0) ack 1015952779 win 5840
 <mss 1460> (DF)
22:43:32.635828 192.168.25.9.14075 > 172.19.64.52.http:
 P 1:590(589) ack 1 win 6144 (DF)
22:43:32.635906 172.19.64.52.http > 192.168.25.9.14075:
 . ack 590 win 6479 (DF)
22:43:32.636758 172.19.64.52.http > 192.168.25.9.14075:
 P 1:217(216) ack 590 win 6479 (DF)
22:43:32.636982 172.19.64.52.http > 192.168.25.9.14075:
 F 217:217(0) ack 590 win 6479 (DF)
22:43:32.639080 192.168.25.9.14075 > 172.19.64.52.http:
 R 590:590(0) ack 217 win 0 (DF)

SNM — ver. 1.4 Network Troubleshooting — slide 95

tcpdump: When reading TCP

• format:

• src ¿ dst: flags data-seqno ack window urgent options

• Flags are some combination of S (SYN), F (FIN), P (PUSH) or R (RST) or a single ‘.’ (no flags).

• The first time tcpdump sees a tcp ’conversation’, it prints the sequence number from the packet.

• On subsequent packets of the conversation, the difference between the current packet’s sequence number and this initial sequence number is printed.

SNM — ver. 1.4 Network Troubleshooting — slide 96

Window

• win ⟨nnn⟩ specifies data window the sending host will accept in

future packets

∘ I.e., the maximum number of bytes

• tcp flow-control:

∘ host reduces this number if congested or overloaded ∘ will sometimes set to 0 to temporarily halt incoming traffic

in this connection

SNM — ver. 1.4 Network Troubleshooting — slide 97

Ethereal King of the Packet Analysers! Available for Linux, Unix, Windows

Ethereal

• Ethereal can read data captured by tcpdump, e.g.,

  $ ethereal -r tcpdump.pca

• or File → Open

• Can capture data itself

• Uses same filter language as tcpdump

SNM — ver. 1.4 Network Troubleshooting — slide 99 SNM — ver. 1.4 Network Troubleshooting — slide 100 SNM — ver. 1.4 Network Troubleshooting — slide 101

You can expand any protocol:

• If we click on the + next to Bootstrap Protocol, we can see the

details of the DHCP Request:

SNM — ver. 1.4 Network Troubleshooting — slide 102 SNM — ver. 1.4 Network Troubleshooting — slide 103

Display Filters

• Note the box at the bottom of Ethereal for display filters

• Select only some of the packets captured for display

• see man ethereal and search for DISPLAY FILTER SYNTAX • Different syntax than the syntax for capture filters

• Example:

• ip.src==172.19.64.52 and ip.dest==172.19.64.57

SNM — ver. 1.4 Network Troubleshooting — slide 104

Tools → Follow TCP Stream

• Can view the contents of an entire tcp stream conversation, in ascii or in hexadecimal.

• Be careful not to invade your customers’ privacy.

• Can use to check if a communications stream is really encrypted

SNM — ver. 1.4 Network Troubleshooting — slide 105

Ntop: monitoring data at a point

• The Ntop program:

∘ listens on a network interface ∘ puts an Ethernet interface into promiscuous mode and ∘ displays statistics through a web interface

• Shows:

∘ percentages of protocols, ∘ which machines generate most traffic ∘ which traffic is purely local, which traffic comes from outside,

which traffic goes from inside to outside of network

SNM — ver. 1.4 Network Troubleshooting — slide 106

Ntop: Installing

• Installation is pretty easy. On my Fedora Core 1 machine:

$ rpmbuild --rebuild ntop-3.0-0.src.rpm
$ sudo rpm -Uhv /home/nicku/RPM/RPMS/i386/ntop-3.0-0.i386.rpm
$ ls -l /etc/ntop.conf*
-rwx------  1 root root 13203 Apr 27 03:47 /etc/ntop.conf.sample
$ sudo cp -a /etc/ntop.conf.sample /etc/ntop.conf
$ sudo emacs /etc/ntop.conf &
# temporarily comment out the line --daemon
$ sudo  /usr/bin/ntop @/etc/ntop.conf -A
$ sudo service ntop start

• Then open the web browser on http://localhost:3000/

SNM — ver. 1.4 Network Troubleshooting — slide 107

Switched Networks Using Ethereal, tcpdump, Ntop in a switched network

Port Monitoring: Switched Networks

Problem:

• a switched network is really a point-to-point network

• You cannot normally capture the unicast traffic from other hosts on a single switch port

• How do you use Ethereal, tcpdump or Ntop to monitor traffic

between a number of hosts?

Solution: many switches support port monitoring, where one port can

monitor all traffic on a specified VLAN Example:

• Cisco 3500XL switches provide the port monitor command:

• port monitor vlan VLAN1

SNM — ver. 1.4 Network Troubleshooting — slide 109

How monitor one machine?

• You are asked to check out a server on a switched network: The switch does not support port monitoring (§109), or you do not have administrative access to the switch: what to do?

• Use a small hub, and use a notebook running the capture software

SNM — ver. 1.4 Network Troubleshooting — slide 110

Are switched networks secure?

• Is all unicast traffic on one port of a switch private?

• No, there are tools (dsniff and Ettercap) freely available to automate arp spoofing and man-in-the-middle attacks, that provide various ways to compromise switch security.

SNM — ver. 1.4 Network Troubleshooting — slide 111

Port Scanning

What is a port scanner?

• Sends packets to various ports on a network device

• Best one available everywhere is nmap

• can identify the OS of the target machine

• Do not port scan arbitrary machines in your company’s network without permission!

• May be interpreted as a cracking attempt

SNM — ver. 1.4 Network Troubleshooting — slide 113

How does nmap identify OS?

• rfcs leave interpretation of some things up to the implementer • rfcs do not specify how should work if get contradictory flags,

strange sequences of inconsistent packets

• Most tcp^{/}ip implementations are not complete

• Every implementation of tcp^{/}ip is different; the “grey areas” are different from one os to another.

• nmap sends “strange” packets to the machine, detects how reacts,

matches this against a file of os fingerprints

SNM — ver. 1.4 Network Troubleshooting — slide 114

Running nmap: Use xnmap

$ sudo -v
$ sudo xnmap &

• Enter the ip address of machine(s) to identify

• select other choices from buttons

• press Start

• xnmap is simply a way to easily generate command line options to

nmap using a graphical interface

SNM — ver. 1.4 Network Troubleshooting — slide 115

Uses of nmap

• Identify the type of a computer that is causing trouble on the network

• Check what network services a computer is really offering ∘ compare with netstat -tua output ∘ A cracked computer may be hiding some services with tro-

janed utilities ∘ nmap can help you discover such services

SNM — ver. 1.4 Network Troubleshooting — slide 116

DNS troubleshooting Troubleshooting DNS Servers

DNS troubleshooting

• Suspect dns when get long timeouts before see any response • ping name, ip address, see if only ip address works

• tools on Linux, Unix:

    ∘ dig, nslookup, host

• tools on Windows:

    ∘ nslookup

SNM — ver. 1.4 Network Troubleshooting — slide 118

DNS: dig

• The people who write the most common name server (Bind) promote dig, deprecate nslookup

• dig output is in form of dns resource records

∘ can copy and paste straight into dns database files

SNM — ver. 1.4 Network Troubleshooting — slide 119

dig: Checking forward DNS lookup

$ dig nicku.org
; <<>> DiG 9.2.1 <<>> nicku.org
;; global options:  printcmd
;; Got answer:
;; ->>HEADER<<- opcode: QUERY, status: NOERROR, id: 23568
;; flags: qr rd ra; QUERY: 1, ANSWER: 1, AUTHORITY: 3, ADDITIONAL: 3
;; QUESTION SECTION:
;nicku.org.            IN      A
;; ANSWER SECTION:
nicku.org.     60      IN      A     202.69.77.139
;; AUTHORITY SECTION:
no-ip.com.              60      IN      NS      nf1.no-ip.com.
no-ip.com.              60      IN      NS      nf2.no-ip.com.
no-ip.com.              60      IN      NS      nf3.no-ip.com.
;; ADDITIONAL SECTION:
nf1.no-ip.com.          60      IN      A       66.185.166.131
nf2.no-ip.com.          60      IN      A       66.185.162.100
nf3.no-ip.com.          60      IN      A       216.66.37.10
;; Query time: 254 msec
;; SERVER: 127.0.0.1#53(127.0.0.1)
;; WHEN: Mon Feb 24 10:55:26 2003
;; MSG SIZE  rcvd: 154

SNM — ver. 1.4 Network Troubleshooting — slide 120

dig: reverse lookup 1

$ dig -x 202.69.77.139
; <<>> DiG 9.2.1 <<>> -x 202.69.77.139
;; global options:  printcmd
;; Got answer:
;; ->>HEADER<<- opcode: QUERY, status: NOERROR, id: 22117
;; flags: qr rd ra; QUERY: 1, ANSWER: 1, AUTHORITY: 2, ADDITIONAL: 0
;; QUESTION SECTION:
;;139.77.69.202.in-addr.arpa.    IN      PTR
;; ANSWER SECTION:
139.77.69.202.in-addr.arpa. 3600 IN     PTR     077-139.onebb.com.
;; AUTHORITY SECTION:
77.69.202.in-addr.arpa. 3600    IN      NS      ns2.onebb.com.
77.69.202.in-addr.arpa. 3600    IN      NS      ns1.onebb.com.
;; Query time: 310 msec
;; SERVER: 172.19.64.52#53(172.19.64.52)
;; WHEN: Mon Feb 24 11:07:04 2003
;; MSG SIZE  rcvd: 111

SNM — ver. 1.4 Network Troubleshooting — slide 121

dig syntax

dig [⟨options⟩] [@⟨server⟩] ⟨name⟩ ⟨type⟩

• main option is -x

• ⟨server⟩ is the name server to query

∘ by default, use first server in /etc/resolv.conf

• ⟨name⟩ is what you want to look up

• ⟨type⟩ can be: any, a, mx, axfr, soa, etc.

• default is to get A record(s)

SNM — ver. 1.4 Network Troubleshooting — slide 122

dig: axfr (Zone Transfer)

• dig can request a complete zone transfer:

$ dig @ns tyict.vtc.edu.hk axfr
; <<>> DiG 9.2.2-P3 <<>> @ns tyict.vtc.edu.hk axfr
;; global options:  printcmd
tyict.vtc.edu.hk.       86400   IN      SOA     ns.tyict.vtc.edu.hk.
 nicku.vtc.edu.hk. 2004031000 3600 1800 604800 600
tyict.vtc.edu.hk.       86400   IN      NS      ns.tyict.vtc.edu.hk.
tyict.vtc.edu.hk.       86400   IN      NS      ns1.tyict.vtc.edu.hk.
tyict.vtc.edu.hk.       86400   IN      NS      dns1.vtc.edu.hk.
tyict.vtc.edu.hk.       86400   IN      NS      dns2.vtc.edu.hk.
000081667.tyict.vtc.edu.hk. 86400 IN    A       172.19.64.92
...

• result can be copied and pasted as a master file in a DNS server

SNM — ver. 1.4 Network Troubleshooting — slide 123

nslookup: an interactive program

$ nslookup
Note:  nslookup is deprecated and may be removed
from future releases.  Consider using the `dig'
or `host' programs instead.  Run nslookup with
the `-sil[ent]' option to prevent this message
from appearing.
> nicku.org
Server:         127.0.0.1
Address:        127.0.0.1#53
Non-authoritative answer:
Name:   nicku.org
Address: 202.69.77.139

SNM — ver. 1.4 Network Troubleshooting — slide 124

nslookup: reverse lookups

> 202.69.77.139
Server:         127.0.0.1
Address:        127.0.0.1#53
Non-authoritative answer:
139.77.69.202.in-addr.arpa      name = 077-139.onebb.com.
Authoritative answers can be found from:
77.69.202.in-addr.arpa  nameserver = ns1.onebb.com.
77.69.202.in-addr.arpa  nameserver = ns2.onebb.com.
ns1.onebb.com   internet address = 202.180.160.1
ns2.onebb.com   internet address = 202.180.161.1
>

SNM — ver. 1.4 Network Troubleshooting — slide 125

Telnet: Troubleshooting Email and Other Protocols

Email: testing with telnet

• Email protocols smtp, pop3 are text

• telnet a good tool to test them

• syntax:

$ telnet ⟨server⟩ ⟨portnumber⟩

• SMTP: port 25

• POP3: port 110

SNM — ver. 1.4 Network Troubleshooting — slide 127

Test the VTC mail server:

$ telnet smtp.vtc.edu.hk 25
Trying 192.168.79.191...
Connected to smtp.vtc.edu.hk (192.168.79.191).
Escape character is '^]'.
220 pandora.vtc.edu.hk ESMTP Mirapoint 3.2.2-GA; Tue, 25 Feb 2003
 11:15:30 +0800 (HKT)
helo nickpc.tyict.vtc.edu.hk
250 pandora.vtc.edu.hk Hello [172.19.32.30], pleased to meet you
mail from: nicku@nicku.org
250 nicku@nicku.org... Sender ok
rcpt to: nicku@nicku.org
250 nicku@nicku.org... Recipient ok
data
354 Enter mail, end with ``.'' on a line by itself
My message body.
.
250 AFF21826 Message accepted for delivery
quit
221 pandora.vtc.edu.hk closing connection
Connection closed by foreign host.

SNM — ver. 1.4 Network Troubleshooting — slide 128

SMTP commands for sending mail

helo identify your computer

mail from specify sender

rcpt to specify receiver

data indicates start of message body

quit terminate session

• Use names, not IP addresses, to specify destination

SNM — ver. 1.4 Network Troubleshooting — slide 129

Testing the VTC pop3 server 1

$ telnet pop.vtc.edu.hk 110
Trying 192.168.79.12...
Connected to pop.vtc.edu.hk (192.168.79.12).
Escape character is '^]'.
+OK carme.vtc.edu.hk POP3 service (iPlanet Messaging Server 5.2
 Patch 1 (built Aug 19 2002))
user nicku
+OK Name is a valid mailbox
pass password
+OK Maildrop ready
stat
+OK 1 673

SNM — ver. 1.4 Network Troubleshooting — slide 130

Testing the pop3 server 2

retr 1
+OK 673 octets
Return-path: <nicku@nicku.org>
Received: from pandora.vtc.edu.hk (pandora.vtc.edu.hk [192.168.79.191])
 by carme.vtc.edu.hk (iPlanet Messaging Server 5.2 Patch 1 (built Aug 19 2002))
 with ESMTP id <0HAU00I35H3HGL@carme.vtc.edu.hk> for nicku@ims-ms-daemon
 (ORCPT nicku@nicku.org); Tue, 25 Feb 2003 11:16:29 +0800 (CST)
Received: from nickpc.tyict.vtc.edu.hk ([172.19.32.30])
        by pandora.vtc.edu.hk (Mirapoint Messaging Server MOS 3.2.2-GA)
        with SMTP id AFF21826; Tue, 25 Feb 2003 11:16:01 +0800 (HKT)
Date: Tue, 25 Feb 2003 11:15:30 +0800 (HKT)
From: Nick Urbanik <nicku@nicku.org>
Message-id: <200302250316.AFF21826@pandora.vtc.edu.hk>
My message body.
.
dele 1
+OK message deleted
quit
+OK
Connection closed by foreign host.

SNM — ver. 1.4 Network Troubleshooting — slide 131

pop3 commands: retrieving mail

• See RFC 1939 for easy-to-read details

• First, must authenticate:

user ⟨username⟩

pass ⟨password⟩

stat shows number of messages and total size in bytes

list list all the message numbers and size in bytes of each message

retr ⟨messagenum⟩ retrieve the message with number ⟨messagenum⟩

dele ⟨messagenum⟩ delete the message with message number ⟨messagenum⟩

quit

SNM — ver. 1.4 Network Troubleshooting — slide 132

telnet: Testing Other Applications

• Many network protocols are text. telnet can be helpful in check-

ing:

∘ IMAP servers:

      $ telnet ⟨hostname⟩ 143

∘ Web servers:

$ telnet ⟨hostname⟩ 80 ∘ Ftp servers:

$ telnet ⟨hostname⟩ 21 ∘ Even ssh (can check version, if responding):

$ telnet ⟨hostname⟩ 22

SNM — ver. 1.4 Network Troubleshooting — slide 133

Conclusion

• Check the simple things first

• Be methodical

• Document what you do

• Become familiar with common tools

• Use the tools to become familiar with your network before troubles strike

• Know what is “normal”

• Getpermission from the boss before using packet sniffing and port

scanners

SNM — ver. 1.4 Network Troubleshooting — slide 134